Show the derivation of the formula.



Cos u/2=±√1+cos u/2.

1
Expert's answer
2022-03-01T11:03:10-0500

cos(u+v) = cos u cos v - sin u sin v

Put v = u, we get

cos (u + u) = cos u cos u - sin u sin u

cos 2u = "cos^2u - sin^2 u"

"cos 2u = cos^2u - (1-cos^2u)"

"cos 2u = 2 cos^2 u - 1"

On rearranging, we get,

"2 cos^2 u = 1 +cos 2u"

"cos ^2u = \frac{1+cos 2u}{2}"

"cos u =\sqrt{(\frac{1+cos2u}{2})}"

put u = u/2

"cos\frac{u}{2} = \sqrt{(\frac{(1+cos u)}{2})}"